A uniform cone of mass 2.0 kg and of height 40 cm, resting on its base, of width 30 cm.​ Calculate:

i) the minimum horizontal pull to the apex which can be applied to raise the edge from the floor. 



1
Expert's answer
2021-01-11T11:41:06-0500

"\tau_1-\tau_2=0\to mg\cdot (\frac{0.3}{2})-F\cdot(0.4)=0"


"F=\frac{mg\cdot (\frac{0.3}{2})}{0.4}=\frac{2\cdot 9.8\cdot (\frac{0.3}{2})}{0.4}=7.35(N)" . Answer